# RadarLLM: Adapting Pretrained Large Language Models for Marine Radar Target Detection with Preference-aware Loss

### v2

@@ III-D2 Theoretical Insight of the Effectiveness of Feature Token Importance
We begin by formalizing the task of selecting informative feature patches for training. At each step $t$, given a candidate set $B_{t}$, our objective is to select a feature patch $(F_{b,k}^{\mathrm{fi}},y_{b}^{k})\in B_{t}$ whose inclusion in the training set $D_{t}$ achieves the maximal reduction in generalization loss on the test set $D_{\mathrm{te}}$. This can be formally formulated as the following optimization problem:
@@
Directly optimizing Eq.(17) is computationally prohibitive, as it entails retraining the model for each candidate feature patch. To derive a tractable surrogate, we reformulate Eq.(17) using Bayes’ theorem:
@@
Assuming that each label is statistically independent of the other pairs in the corpus when conditioned on its own feature vector, the joint term in the numerator can be expressed as:
@@
Substituting Eq.(19) into Eq.(18) gives:
@@
Eq.(21) indicates that feature patches with larger loss differences contribute more significantly to improve the model performance on test set; thus, the loss difference can be interpreted as the learning value of each feature patch. However, since accessing the test set for retraining on $D_{t}\cup D_{\mathrm{test}}$ is prohibited and computationally impractical, we approximate the second term using a lightweight reference model trained solely on validation set $D_{\mathrm{val}}$. Consequently, the practical token weight is derived as shown in Eq.(14).
@@ III-D3 Comparison with Existing Loss Designs
III-E False Alarm Controllable Classification Head Retraining
As shown in Fig. 1, the LLM-based model outputs $[B,K,2]$ token-level predictions, which lack a unified detection representation. To achieve consistent decision-making, we retrain a classifier using the learned feature $F^{\mathrm{LN}}$. An autoencoder-based classification head extracts informative features from this high-dimensional input for robust detection. It comprises: (1) an encoder with convolutional and fully connected layers producing compact discriminative features; (2) a symmetric decoder ensuring feature consistency through reconstruction; and (3) a classification head operating on the latent vector. The autoencoder is optimized by a dual-objective loss combining reconstruction and classification terms:
@@ III-E False Alarm Controllable Classification Head Retraining
During inference, false alarms are controlled via threshold-setting using sorted softmax outputs of clutter samples. The detection threshold $\eta$ is then determined based on the desired false alarm rate $P_{fa}^{\text{d}}$ as follows:
@@
where $O_{\mathrm{sorted}}(x)$ is the $x$-th largest output and $N_{\text{clutter}}$ is the total number of test clutter samples.
@@ IV Experiments
In this section, we first describe the experimental setup, followed by a comprehensive evaluation of RadarLLM’s detection performance under two settings: sufficient training data and constrained training data scenarios.
IV-A Experimental Setup
IV-A1 Dataset
@@
IV-A2 Baselines
To validate the superiority of RadarLLM, eight deep learning-based marine radar target detection methods are implemented as baselines. Among them, RNN, Bi-LSTM, GRU, Transformer, PatchTST, and ResNet18 represent lightweight models that utilize sequence features; OFA uses a large language model backbone to process sequence features; and ADN18 utilizes time-frequency maps and enhanced CNNs for detection.
IV-A3 Evaluation Metrics
IV-A4 Implementation Details
@@
Learning rate for LLM fine-tuning 0.0001
@@
IV-B Comparison of Detection Performance, Speed, and Network Parameters Under Sufficient Training Samples
@@
TABLE IV: Detection performance of different methods when the false alarm rate is 0.005. The best and second-best detection results are highlighted in red and blue, respectively.
Detector $H$ HV VH $V$ AVG
@@
RadarLLM 82.51 84.81 84.61 75.90 81.96
TABLE V: Network parameters (training parameters/total parameters) and interference time of different methods.
Detector Average interference time on one dataset(s) Total network parameters(M) Training network parameters(M)
@@
Table IV reports the detection rates achieved by RadarLLM and the eight baseline detectors introduced in Section IV-A2. All results are obtained with a fixed false alarm rate of 0.005 on the eight IPIX datasets. Compared to the best competing sequence feature detector, RadarLLM improves the detection rate by 7.60%, 3.85%, 2.30%, and 1.93% for the HH, HV, VH, and VV polarizations, respectively. Figure IV presents the complete ROC curves; it is evident that RadarLLM maintains the highest detection probability across the full range of false alarm rates. We also compare RadarLLM with ADN18, which relies on high-dimensional time–frequency maps and an enhanced CNN backbone. Despite using only compact sequence features, RadarLLM still delivers a 1.14% gain in the average detection rate.
@@
Furthermore, we compare RadarLLM with other baseline models in terms of training complexity, total network parameters, and inference cost to evaluate its practicality for real-world deployment. The comparison results are summarized in Table V. Although RadarLLM has the highest total network parameters, the number of trainable parameters is comparable to those of other models, as the majority of its weights are frozen during fine-tuning. In particular, RadarLLM achieves an exceptionally low average inference latency. Specifically, it requires only 5.6561 seconds to process more than 9,000 test samples, which is 28.44 times faster than the ADN18 method. This efficiency is primarily attributed to our patching strategy, which substantially reduces the number of feature tokens, and consequently, the computational overhead. Moreover, the intrinsic inference acceleration of the GPT architecture further enhances this performance gain. Together, these advantages make RadarLLM a promising solution for real-time marine target detection.
IV-C Comparison of Detection Performance Under Limited Training Samples With Baselines
@@
Due to the inherent sparsity of radar targets in real-world scenarios, obtaining high-quality labeled samples for radar target detection presents a substantial challenge. Considering this issue, we conducted experiments under constrained training sample conditions to assess the effectiveness of RadarLLM. The samples are divided into three groups: (1) a training set using the first 10% of the observation time for both the target and clutter cells, (2) a validation set covering 10% to 15% of the observation time, and (3) a test set containing the remaining samples.
@@
In Table VI, we report the detection performance of RadarLLM and several baseline methods under scenarios with limited training data. Compared with the experimental setup described in Section IV-A1, the advantages of RadarLLM become more pronounced. Specifically, RadarLLM achieves a substantial increase of 19.66% in the average detection rate over the previously proposed sequence-feature based detector [14]. This marked improvement can be mainly attributed to three key factors:
  • First, leveraging pre-trained knowledge embedded in the large language model (LLM) provides RadarLLM with a beneficial parameter initialization, facilitating efficient optimization even when training data is limited. This prior knowledge significantly reduces the data dependency typical of conventional training strategies.
@@
  • Second, the substantial parameter scale of RadarLLM inherently grants robustness to distributional variations, enabling superior generalization compared to smaller-scale neural networks. Its extensive capacity allows it to capture complex data structures and efficiently handle variability in different radar detection scenarios.
@@
  • Third, the integration of a preference-aware loss function significantly improves RadarLLM detection performance, especially in challenging scenarios characterized by low signal-to-clutter ratios (SCR).
TABLE VI: Detection performance under limited training samples of different methods when the false alarm rate is 0.005. The best and second-best detection results are highlighted in red and blue, respectively.
Detector $H$ HV VH $V$ AVG
@@
RadarLLM 79.26 82.31 81.51 72.54 78.91
(1). Data1 HH
(2). Data1 HV
(3). Data1 VH
(4). Data1 VV
(5). Data2 HH
(6). Data2 HV
(7). Data2 VH
(8). Data2 VV
(9). Data3 HH
(10). Data3 HV
(11). Data3 VH
(12). Data3 VV
Figure 7: Comparison of cross-entropy loss and preference-aware loss on low SCR datasets. We compare detection rates across different datasets and epochs, highlighting the improvements achieved by the preference-aware loss function.
@@ IV-D Ablation Study
IV-D1 Ablation Study on Backbone Architectures and Fine-tuning Strategies
To validate the effectiveness of the backbone choice, the ablation study is conducted in this section by removing or replacing the pre-trained LLM:
@@ IV-D1 Ablation Study on Backbone Architectures and Fine-tuning Strategies
    RadarLLM (0): In this variant, the pre-trained LLM is entirely removed, while all other components of the framework are kept unchanged.
@@
    RadarLLM (T): In this variant, the pre-trained LLM is replaced with a randomly initialized transformer block.
@@
    RadarLLM (R): In this variant, the network architecture and fine-tuning procedure of the original LLM are maintained, but its pre-trained parameters are replaced with random initialization.
@@
In Table VII, we show the results of the ablation study on backbone choice. It can be seen that the removal or exchange of pre-trained LLM results in notable performance degradation, indicating the necessity of pretrained LLM for high detection performance.
@@
TABLE VII: Results of ablation study on the backbone choice.
$H$ HV VH $V$ AVG
RadarLLM 82.51 84.81 84.61 75.90 81.96
RadarLLM (0) 77.51 79.38 80.35 69.97 76.80
RadarLLM (T) 78.25 79.77 81.16 72.13 77.83
RadarLLM (R) 75.19 75.44 75.81 65.15 72.90
Moreover, to verify the effectiveness of the proposed fine-tuning strategy, we conducted an ablation study as summarized in Table VIII, and the corresponding results are presented in Table IX. The results demonstrate that jointly fine-tuning the LoRA modules and Layer Normalization parameters achieves the highest overall performance. In contrast, fine-tuning only the Layer Normalization parameters, the LoRA modules, or performing full fine-tuning results in moderate performance degradation. These findings indicate that the proposed lightweight fine-tuning strategy effectively balances model adaptability and computational efficiency while preserving the universal knowledge embedded in the LLM backbone.
TABLE VIII: Ablation Study on Fine-tuning Strategies settings.
@@
LoRA Layer Norm Other Params
RadarLLM $\checkmark$ $\checkmark$
RadarLLM (F) $\checkmark$ $\checkmark$
RadarLLM (LoRA) $\checkmark$
RadarLLM (LN) $\checkmark$
TABLE IX: Results of ablation study on the finetuning strategy.
$H$ HV VH $V$ AVG
RadarLLM 82.51 84.81 84.61 75.90 81.96
RadarLLM (F) 81.69 83.45 84.41 74.07 80.91
RadarLLM (LoRA) 82.20 84.45 84.56 75.53 81.69
RadarLLM (LN) 82.33 83.73 83.80 75.12 81.25
IV-D2 Ablation Study on the Loss Function

### v3

@@ III-D2 Theoretical Insight of the Effectiveness of Feature Token Importance
A. Problem Definition: Inspired by [26], we begin by formalizing an optimization problem of selecting useful feature patches for model fine-tuning. At each training step $t$, given a candidate set $B_{t}$, the objective is to select a feature patch $(F_{b,k}^{\mathrm{fi}},y_{b}^{k})\in B_{t}$ whose inclusion in the training set $D_{t}$ maximally reduces the generalization loss on the unseen test set $D_{\mathrm{te}}$. This can be expressed as the following optimization problem:
@@
However, directly optimizing Eq.(16) is computationally prohibitive, as it requires retraining the model for each candidate patch. To enable a practical surrogate, we reformulate this objective using probabilistic decomposition.
B. Derivation of the Surrogate Objective: Applying Bayes’ theorem, the term in Eq.(16) can be decomposed as:
@@
Assuming that each label is statistically independent of all other labels in the corpus when conditioned on its corresponding feature vector, the joint probability term in the numerator can be expressed as follows:
@@
Substituting Eq.(18) into Eq.(17) and taking the logarithm yields:
@@
As exact Bayesian inference is intractable in neural networks, we fit the models with SGD instead. Here, $\mathcal{L}[y_{b}^{k}\mid F_{b,k}^{\mathrm{fi}};D_{t}]$ denotes the training loss on the patch using the current model trained on $D_{t}$, whereas $\mathcal{L}\left[y_{b}^{k}\mid F_{b,k}^{\mathrm{fi}};D_{\mathrm{t}},D_{\mathrm{test}}\right]$ represents the excepted irreducible holdout loss obtained if the model were trained on both the training set and the test set.
C. Conclusion: Eq.(20) reveals that the gap between the training loss and the irreducible holdout loss quantifies the reducible component of the generalization error of each feature patch. Consequently, this loss difference can be rigorously interpreted as the learning value of each feature patch, representing how much its inclusion in model training is expected to improve performance on unseen data.
However, accessing the test set for training on $D_{t}\cup D_{\mathrm{test}}$ is both prohibited and computationally impractical. To address this limitation, the validation set is adopted as a statistically sound and feasible proxy for the unseen data distribution. Accordingly, we approximate $\mathcal{L}\left[y_{b}^{k}\mid F_{b,k}^{\mathrm{fi}};D_{\mathrm{t}},D_{\mathrm{test}}\right]$ using the loss computed by a lightweight reference model trained on the validation set $D_{\mathrm{val}}$ [26]. Consequently, the resulting token weight derived in Eq. (14) provides a practical and theoretically grounded measure of each feature patch’s learning value.
@@ III-D3 Comparison with Existing Loss Designs
III-E Autoencoder-based Binary Classification Head Retraining
As shown in Fig. 1 and Section III-A, the PLM-based model outputs $[B,K,2]$ token-level predictions, which lack a unified detection representation. To achieve consistent decision-making, we retrain a binary classifier head using the learned representation $F^{\mathrm{LN}}$. An autoencoder-based biary classification head extracts informative features from this high-dimensional input for robust detection. It comprises: (1) an encoder with convolutional and fully connected layers producing compact discriminative representation; (2) a symmetric decoder ensuring feature consistency through reconstruction; and (3) a classification head operating on the latent vector. The autoencoder is optimized by a dual-objective loss combining reconstruction and classification terms:
@@ III-E Autoencoder-based Binary Classification Head Retraining
During inference, false alarm rate is controlled via threshold-setting using sorted softmax outputs of clutter samples. The detection threshold $\eta$ is then determined based on the desired false alarm rate $P_{fa}^{\text{d}}$ as follows:
@@
where $O_{\mathrm{sorted}}(x)$ is the $x$-th largest output and $N_{\text{clutter}}$ is the total number of test clutter samples. The overall procedure of the RadarPLM is presented in Algorithm 1.
Algorithm 1 RadarPLM Training Protocol
1:
2:Radar echo signals $\{x(n)\}_{n=1}^{N}$ and Target labels $\{y_{b}\}_{b=1}^{B}$
3:Pre-trained LLM parameters $\theta_{\text{LLM}}$
4:Stage 1: Sequence Feature Extraction
  1. 1.
  2. 2.
    Patch the five features into feature tokens, see Fig. 3.
5:Stage 2: Reference Model Training
  1. 1.
  2. 2.
6:Stage 3: Fine-tune PLM for Target Detection
  1. 1.
  2. 2.
    Fine-tune PLM for target detection via preference-aware loss function, see Eq.(14) and Eq.(15).
7:Stage 4: Binary Classification Head Retraining
  1. 1.
    Train autoencoder-based binary classification head based on the weighted sum of reconstruction loss and CE loss, see Eq.(21), Eq.(22), and Eq.(23).
  2. 2.
    Adjust detection threshold $\eta$ for controllable FAR, see Eq.(24).
8:
9:Fine-tuned RadarLLM model with controllable FAR.
@@ IV Experiments
Learning rate for PLM fine-tuning 0.0001
@@
In this section, we first describe the experimental setup, followed by a comprehensive evaluation of RadarPLM’s detection performance under two settings: sufficient training data and limited training data scenarios.
IV-A Experimental Setup
IV-A1 Dataset
IV-A2 Baselines
To validate the superiority of RadarPLM, eight deep learning-based marine radar target detection methods are implemented as baselines. Among them, RNN, Bi-LSTM, GRU, Transformer, PatchTST, and ResNet18 represent lightweight models that process sequence features; OFA uses a PLM backbone to process sequence features; and ADN18 utilizes time-frequency maps and enhanced CNNs for detection.
IV-A3 Evaluation Metrics
IV-A4 Implementation Details
IV-B Performance Comparison on Sufficient Training Samples
@@
TABLE IV: Comparison of detection rates (%) on sufficient training samples.
Method $H$ HV VH $V$ AVG
@@
RadarPLM 82.51 84.81 84.61 75.90 81.96
(All results are evaluated at a false alarm rate ($P_{fa}$) of 0.005. The best and second-best results are highlighted in red and blue, respectively.)
TABLE V: Network parameters (training parameters / total parameters) and average inference time of different methods.
Detector Avg. inference time (s) Total Params (M) Trainable Params (M)
@@
Table IV reports the detection rates achieved by RadarLLM and the eight baseline detectors introduced in Section IV-A2. All results are obtained with a fixed false alarm rate of 0.005 on the eight IPIX datasets. Compared to the best competing sequence feature detector, RadarLLM improves the detection rate by 7.60%, 3.85%, 2.30%, and 1.93% for the HH, HV, VH, and VV polarizations, respectively. Figure LABEL:DT presents the complete ROC curves; it is evident that RadarLLM maintains the highest detection probability across the full range of false alarm rates. We also compare RadarLLM with ADN18, which relies on high-dimensional time–frequency maps and an enhanced CNN backbone. Despite using only compact sequence features, RadarLLM still delivers a 1.14% gain in the average detection rate.
@@
Furthermore, we compare RadarPLM with other baseline models in terms of training complexity, total network parameters, and inference cost to evaluate its practicality for real-world deployment. The comparison results are summarized in Table V. Although RadarPLM has the highest total network parameters, the number of trainable parameters is comparable to those of other models, as the majority of its weights are frozen during fine-tuning. In particular, RadarPLM achieves an exceptionally low average inference latency. Specifically, it requires only 5.6561 seconds to process more than 9,000 test samples, which is 28.44 times faster than the ADN18 method. This efficiency is primarily attributed to our patching strategy, which substantially reduces the number of feature tokens, and consequently, the computational overhead. Moreover, the intrinsic inference acceleration of the GPT architecture further enhances this performance gain. Together, these advantages make RadarPLM a promising solution for real-time marine target detection.
IV-C Performance Comparison on Limited Training Samples
@@
Due to the inherent sparsity of radar targets in real-world scenarios, obtaining high-quality labeled samples for radar target detection presents a substantial challenge. Considering this issue, we conducted experiments under constrained training sample conditions to assess the effectiveness of RadarPLM. The samples are divided into three groups: (1) a training set using the first 10% of the observation time for both the target and clutter cells, (2) a validation set covering 10% to 15% of the observation time, and (3) a test set containing the remaining samples.
@@
In Table VI, we report the detection performance of RadarPLM and several baseline methods under scenarios with limited training data. Compared with the experimental setup described in Section IV-A1, the advantages of RadarPLM become more pronounced. Specifically, RadarPLM achieves a substantial increase of 19.66% in the average detection rate over the previously proposed sequence-feature based detector [14]. This marked improvement can be mainly attributed to three key factors:
  • First, The embedded knowledge from pre-trained parameters endows RadarPLM with a highly advantageous parameter initialization. This acts as a powerful springboard for optimization, enabling effective convergence even in data-limited regimes and drastically mitigating the data-hungry nature of conventional training paradigms.
@@
  • Second, the substantial parameter scale of RadarPLM inherently grants robustness to distributional variations, enabling superior generalization compared to smaller-scale neural networks. Its extensive capacity allows it to capture complex data structures and efficiently handle variability in different radar detection scenarios.
@@
  • Third, the integration of a preference-aware loss function significantly improves RadarPLM detection performance, especially in challenging scenarios characterized by low SCR.
TABLE VI: Comparison of detection rates (%) on limited training samples.
Method $H$ HV VH $V$ AVG
@@
RadarPLM 79.26 82.31 81.51 72.54 78.91
(All results are evaluated at a false alarm rate ($P_{fa}$) of 0.005. The best and second-best results are highlighted in red and blue, respectively.)
(1). Data1
(2). Data2
(3). Data3
Figure 7: Comparison of detection rates between the proposed preference-aware loss and the conventional cross-entropy loss. We compare detection rates across different datasets and epochs, highlighting the improvements achieved by the preference-aware loss function.
@@ IV-D Ablation Study
IV-D1 Backbone Architectures
To validate the effectiveness of the backbone choice, the ablation study is conducted in this section by removing or replacing the PLM under sufficient training samples setting:
@@ IV-D1 Backbone Architectures
    RadarPLM (0): In this variant, the PLM backbone is entirely removed, while all other components of the framework are kept unchanged.
@@
    RadarPLM (T): In this variant, the PLM backbone is replaced with a randomly initialized transformer block.
@@
    RadarPLM (R): In this variant, the network architecture and fine-tuning procedure are maintained, but pre-trained parameters of PLM are replaced with random initialization.
@@
In Table VII, we show the results of the ablation study on backbone choice. It can be seen that the removal or exchange of PLM backbone results in notable performance degradation, indicating the necessity of PLM for high detection performance.
@@
TABLE VII: Results of ablation study on the backbone choice and fine-tuning strategies. RadarPLM(0)–RadarPLM(R) denote backbone ablation results, while RadarPLM(F)–RadarPLM(LN) denote fine-tuning strategy ablation results.
Model $H$ HV VH $V$ AVG
RadarPLM 82.51 84.81 84.61 75.90 81.96
Backbone choice ablation
RadarPLM (0) 77.51 79.38 80.35 69.97 76.80
RadarPLM (T) 78.25 79.77 81.16 72.13 77.83
RadarPLM (R) 75.19 75.44 75.81 65.15 72.90
Fine-tuning strategy ablation
RadarPLM (F) 81.69 83.45 84.41 74.07 80.91
RadarPLM (LoRA) 82.20 84.45 84.56 75.53 81.69
RadarPLM (LN) 82.33 83.73 83.80 75.12 81.25
IV-D2 Fine-tuning Strategies
To verify the effectiveness of the proposed fine-tuning strategy, we conducted an ablation study under the setting of sufficient training samples, as summarized in Table VIII, and the corresponding results are presented in Table VII. In this experiment, four fine-tuning configurations were evaluated: (1) RadarPLM: jointly fine-tuning the LoRA modules and Layer Normalization (LN) parameters, (2) RadarPLM (LoRA): fine-tuning only the LoRA modules, (3) RadarPLM (LN): fine-tuning only the LN parameters, and (4) RadarPLM (F): full fine-tuning of all model parameters. As shown in the results, the joint fine-tuning of LoRA and LN achieves notably higher performance, while maintaining a relatively shorter training cost (0.116 h/100 epochs) than full fine-tuning (0.148 h/100 epochs). This demonstrates that the proposed design not only enhances model adaptability, but also significantly reduces computational cost.
TABLE VIII: Ablation study on fine-tuning strategy configurations and their corresponding training time (measured over 100 training epochs).
@@ IV-D2 Fine-tuning Strategies
Model LoRA Layer Norm Other Params Training Time (hour/100 epochs)
RadarPLM $\checkmark$ $\checkmark$ 0.116
RadarPLM (F) $\checkmark$ $\checkmark$ 0.148
RadarPLM (LoRA) $\checkmark$ 0.116
RadarPLM (LN) $\checkmark$ 0.112
IV-D3 Loss Function